# Convergence and error control of consistent PINNs for elliptic PDEs

### v1

@@ Abstract
We provide an a priori analysis of a certain class of numerical methods, commonly referred to as collocation methods, for solving elliptic boundary value problems. They begin with information in the form of point values of the right side $f$ of such equations and point values of the boundary function $g$ and utilize only this information to numerically approximate the solution $u$ of the Partial Differential Equation (PDE). For such a method to provide an approximation to $u$ with guaranteed error bounds, additional assumptions on $f$ and $g$, called model class assumptions, are needed. We determine the best error (in the energy norm) of approximating $u$, in terms of the number of point samples $m$, under all Besov class model assumptions for the right hand side $f$ and boundary $g$.
We then turn to the study of numerical procedures and asks whether a proposed numerical procedure (nearly) achieves the optimal recovery error. We analyze numerical methods which generate the numerical approximation to $u$ by minimizing a specified data driven loss function over a set $\Sigma$ which is either a finite dimensional linear space, or more generally, a finite dimensional manifold. We show that the success of such a procedure depends critically on choosing a correct data driven loss function that is consistent with the PDE and provides sharp error control. Based on this analysis a loss function ${\cal L}^{*}$ is proposed.
We also address the recent methods of Physics Informed Neural Networks (PINNs). Minimization of the new loss ${\cal L}^{*}$ over neural network spaces $\Sigma$ is referred to as consistent PINNs (CPINNs). We prove that CPINNs provides an optimal recovery of the solution $u$, provided that the optimization problem can be numerically executed and $\Sigma$ has sufficient approximation capabilities. Finally, numerical examples illustrating the benefits of the CPINNs are given.
@@ 1 Introduction
where $\partial\Omega$ is the boundary of $\Omega$. In order to prove the existence and uniqueness of a solution to (1), one needs to impose conditions on $f,g,a$, and $\Omega$. A standard set of assumptions that guarantee the existence and uniqueness of a solution (1) is the following: A1: $\Omega$ is a Lipschitz (graph) domain, i.e., $\Omega$ is an open connected set in $\mathbb{R}^{d}$ with a Lipschitz boundary in the sense described in [57]; A2: $f\in H^{-1}(\Omega)$; A3: $g\in H^{1/2}(\partial\Omega)$; A4: the diffusion coefficient $a$ satisfies the coercivity condition
$0<r\leq a({\bf x})\leq M,\quad{\bf x}\in\Omega,$ (1.2)
for some constants $r,M>0$. Here the spaces $H^{s}(\Omega)$, $s\in\mathbb{R}$, are the Sobolev spaces of order $s$ in $L_{2}(\Omega)$.
@@
where the constants $c,C$ depend on $\Omega$, $r$ and $M$  [^1] Throughout this paper, we use the notation $c,C,c_{1},C_{1}$, etc, to denote constants. Unless they are absolute constants, we indicate the quantities on which they depend when they occur. The value of these constants can change at each occurrence. We also use the notation $A\asymp B$ to denote that $c_{1}A\leq B\leq c_{2}A$ with $c_{1},c_{2}>0$.. Therefore, if we consider the theoretical loss function
@@
then minimizing this loss over the whole of $H^{1}(\Omega)$ has $u$ as its unique solution.
In order to make our presentation as clear as possible, we want to avoid the technicalities of handling general domains and general diffusion coefficients $a$. Accordingly, we restrict our presentation to the case $\Omega=(0,1)^{d}$ and $a\equiv 1$. This will allow us to concentrate on the novelties of PINNs and alternative collocation methods. The treatment of more general settings for the diffusion $a$ and domains $\Omega$ is left to future works. We also want to point out that assumptions A2-A3 could be replaced with other smoothness assumptions on $f$ and $g$, respectively, which will result in a theory similar to the one described in this paper.
@@ 1.1 Numerical methods for solving (1)
A numerical method assumes that $f,g$, and $\Omega$ are known and suggests a procedure for solving (1) based on that information. The numerical procedure generates a function $\hat{u}$ which is an approximation to $u$. The efficiency of the numerical method is then evaluated in the following sense. One chooses a norm $\|\cdot\|_{X}$ in which to measure performance and then seeks to quantify how fast the error $\|u-\hat{u}\|_{X}$ between the true solution $u$ and the output $\hat{u}$ of the numerical procedure tends to zero in terms of the complexity of the numerical procedure. The classical norm $\|\cdot\|_{X}$ used to measure error is the energy norm which corresponds to choosing $X=H^{1}(\Omega)$. Other function norms sometimes used to measure error correspond to $X=L_{p}(\Omega)$, $1\leq p\leq\infty$. We restrict our analysis of convergence and rates of convergence to the case $X=H^{1}(\Omega)$ in going further.
An important question is how one can compare the performance of different numerical methods for solving (1) in order to give a fair competition between all possible numerical methods. This is typically done by assigning some form of complexity to the numerical method. In the case of FEMs and AFEMs, this complexity can be measured in terms of the number $n$ of machine operations used to compute $\hat{u}_{n}=\hat{u}$ and there are theorems that give a priori bounds for the error $\|u-\hat{u}_{n}\|_{H^{1}(\Omega)}$ under additional assumptions on the smoothness of $f$ and $g$ (see e.g. [5]). Some alternatives to measuring the complexity $n$ in terms of machine operations are the following. If the numerical approximation $\hat{u}$ to $u$ is an element from a linear $n$-dimensional space $V_{n}$, then the associated complexity is typically assigned to be the dimension $n$. This may not be directly converted to an appropriate number of machine operations because numerical stability becomes an issue. If the approximation $\hat{u}$ comes from a nonlinear manifold ${\cal M}_{n}$, then one can use the number of parameters $n$ determining the manifold as a complexity measure. However, in the latter case, one has to impose extra conditions on the manifold (or the numerical procedure) to prevent the use of space filling manifolds and thus avoid unstable numerical methods, see [9]. In the absence of such restrictions, the numerical method may require an inordinate amount of computational resources to achieve a desired accuracy even in the case when the dimension $n$ is small.
@@
A second item in collocation methods is how one numerically uses the $m$ pieces of information provided to create a numerical approximation to $u$. If $\Sigma_{n}$ is the underlying space used to create $\hat{u}$ then the accuracy of the numerical recovery will depend on $n$ and the choice of $\Sigma_{n}$. In other words, $\hat{u}$ can be viewed as $\hat{u}=\hat{u}_{n,m}$. If one fixes $m$, then one can study how fast the accuracy of approximation $\|u-\hat{u}_{n,m}\|$ tends to the optimal recovery rate as $n\to\infty$. This issue is discussed in §8.
@@ 1.2 Collocation methods and PINNs
$\hat{u}\,\in\,\mathop{\rm argmin}_{S\in\Sigma_{n}}{\cal L}_{0}(S),$ (1.6)
@@
${\cal L}_{0}(S):=\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta S({\bf x}_{i})+f({\bf x}_{i})]^{2}+\frac{\lambda}{\overline{m}}\sum_{i=1}^{\overline{m}}[S({\bf z}_{i})-g({\bf z}_{i})]^{2},$ (1.7)
with $\lambda$ being a tuning parameter. For simplicity, throughout the paper we will set $\lambda=1$ and consider the loss
${\cal L}(S):=\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta S({\bf x}_{i})+f({\bf x}_{i})]^{2}\right]^{1/2}+\left[\frac{1}{\overline{m}}\sum_{i=1}^{\overline{m}}[S({\bf z}_{i})-g({\bf z}_{i})]^{2}\right]^{1/2}$ (1.8)
in our analysis. Note that this is equivalent to the square root of the typical PINNs loss with $\lambda=1$. The appropriateness of the loss ${\cal L}_{0}$ and how $\widetilde{m},\overline{m},\lambda$ and the data sites $({\cal X},{\cal Z})$ should be chosen are major issues and one of the focal points of this paper.
The PINNs numerical procedures fall into the broad class of collocation methods for solving PDEs, i.e. numerical methods that use only values of $f$ and $g$ at specified data sites ${\cal X}$ and ${\cal Z}$. Such methods were studied in the last century (see e.g. [1, 34, 31]) but became less popular with the advent of FEMs. The novelty of PINNs is to use neural networks instead of polynomials or splines to build the approximation $\hat{u}$ from the given data.
@@ 1.3 Overview of this paper
The idea of using neural networks to solve PDEs goes back to the last century [20, 39], and has recently become extremely popular with the introduction of physics-informed neural networks (PINNs) [50]. Despite the increasing usage of PINNs for numerically solving a wide range of PDEs (see for example, [36, 6, 10, 6, 43] and the references therein), a satisfactory analysis of the convergence and performance of these methods has not been given. Some partial progress towards this goal has recently been made, see for instance [24, 45, 53, 54, 65].
@@
For example, in [53], convergence is proved in the $C(\overline{\Omega})$-norm for a modified physics-informed neural network for elliptic PDEs under the assumption that the right hand side $f$ and boundary data $g$ are Hölder continuous. Under the additional assumption that the outputs of the neural network satisfy the boundary conditions exactly, convergence is obtained in $H^{1}(\Omega)$. In this situation, the boundary values are not enforced via the loss function and must instead be implemented through the neural network architecture. We remark that although methods for exactly enforcing boundary conditions with neural networks have been proposed in [39, 40, 3], this approach appears unable to rigorously handle arbitrary boundary conditions on general domains.
@@
Further, in [54] and [65], an abstract framework for analyzing methods based upon residual minimization, such as PINNs, is presented. This framework has been used in a variety of follow up works to analyze PINNs (see [2, 58, 22] for instance). However, the current analysis is, to the best of our knowledge, not able to obtain convergence rates with respect to the number of collocation points when a fixed (as opposed to random) set of collocation points are used.
An analysis taking into account the rate of convergence with respect to the number of collocation points has been obtained in [45]. It applies to the heat equation (instead of elliptic PDEs) under the assumption that the initial data lies in the Sobolev space $H^{s}(\Omega)$ for sufficiently large $s$ to ensure that point values are well-defined.
@@
However, so far, none of the existing literature treats the problem of obtaining convergence rates in terms of both the number of collocation points and the network size for elliptic PDEs under general Besov regularity assumptions on the right hand side $f$ and boundary value $g$. A proper a priori analysis that addresses this problem would answer the following questions: Q1: Given the model class assumptions on $f$ and $g$, which query sites should be used and how large would we have to take $\widetilde{m}$ and $\overline{m}$ to be able to reach a target accuracy $\varepsilon$ for the error $\|u-\hat{u}\|_{H^{1}(\Omega)}$? Q2: Given a budget of $m$ queries and given model class assumption on $f$ and $g$, what is the smallest error that can be achieved in the recovery of $u$? This is called the optimal recovery accuracy. Q3: If we use a NN space $\Sigma_{n}$ with $n$ parameters to build the approximation $\hat{u}$ to $u$, then how large do we need to choose $n$ to achieve near optimal accuracy? Q4: If we use minimization of a data driven loss function to generate the numerical approximation $\hat{u}$, then what form should the loss function take so that small values of this loss function guaranees small values of the solution approximation error, up to the optimal recovery rate ?
The goal of the present paper is to answer the above questions for all Besov space model classes and thereby establish such an a priori analysis of collocation methods and in particular PINNs. In contrast to previous works, we also consider the error in the $H^{1}(\Omega)$-norm, and treat the situation where the boundary values are enforced through the loss function. As already noted, in order to simplify our presentation, we will only consider the case of Laplace’s equation (1) with $a\equiv 1$ on $\Omega$ and $\Omega=(0,1)^{d}$. We leave the problem of handling more general settings to future work. We remark that an analysis of other methodologies for solving (1) using neural networks, such as the deep Ritz method [60] and finite neuron method [61], can be found in [42, 41, 56, 47, 23, 21].
@@
As we have already mentioned, any a priori convergence analysis of collocation methods requires assumptions on the functions $f$ and $g$ in the form that $f\in{\cal F}$, $g\in{\cal G}$, where ${\cal F}$ and ${\cal G}$ are compact subsets of the space of continuous functions on the domains $\overline{\Omega}$ and $\partial\Omega$, respectively. Such compact sets are typically described by smoothness conditions. In this paper, we measure smoothness by membership in Besov classes. For this reason, we begin in §2 with the definition of Besov spaces and the properties we will need in the sequel. That section will also discuss how well functions in Besov classes can be approximated (in various norms) by piecewise polynomials. All of the results in that section are known and are therefore presented in a summary form. However, for completeness of the paper, and for the specific presentation of these results, we give proofs of the results we need in the appendix.
@@
We use membership in Besov spaces $B_{q}^{s}(L_{p}(\Omega))$ as the smoothness we impose on $f$ and $g$. Once such assumptions are placed on $f$ and $g$, we answer questions Q1-Q2 in §3. Questions of this type are commonly referred to as optimal recovery (OR) of a function $u$ from data. Since our data is given by point samples, this is also known as optimal sampling. Our general setting is not usually addressed in the OR literature because our data is not point evaluations of the target function $u$, but rather of $f$ and $g$. Nevertheless, we consider the results of §3 to be for the most part known in the sense that we are simply piecing together various known results and techniques such as those in [12, 13, 38, 49]. The paper [59], which studies optimal sampling with respect the negative Sobolev norm $H^{-1}(\Omega)$ on Lipschitz domains $\Omega$, comes closest to presenting the optimal recovery results in the form we need. However, note that the cases $10<p,q<$ 1 for all $d$, or $p=1$, $0<q\leq\infty$, $d=2$ are not covered in [59]. Notably, we observe that the optimal sampling rates for Besov spaces in $H^{-1}(\Omega)$ coincide with the optimal sampling rates in a certain space $L_{\gamma}(\Omega)$, determined by the Sobolev embedding (see Theorem 3.1 (iv)).
@@
The results of §3 do not give a numerical method for finding $\hat{u}$ from the given data that provides the error bound $\|u-\hat{u}\|_{X}\leq\varepsilon$. They simply establish the theoretical benchmark for the optimal accuracy of any numerical procedure for the recovery of $u$ based on the point samples of $f$ and $g$. This theoretical analysis does not involve NNs or any other proposed collocation method. NNs and PINNs only enter the picture when one wants to transform the theoretical analysis into a numerical procedure that utilizes NNs to find an approximant to $u$.
@@
We next discuss in §4 the use of classical numerical methods such as Finite Element Methods (FEMs) for optimal recovery. In §5, we turn to the question of using optimization to solve the PDE. We observe that minimizing the theoretical loss ${\cal L}_{T}$ over a space $\Sigma_{n}$ will approximate well $u$, provided $\Sigma_{n}$ has sufficient approximation power. However, this is not a numerical algorithm per se because it does not provide a numerical recipe for evaluating ${\cal L}_{T}(S)$ when $S\in\Sigma_{n}$.
@@
The remainder of this paper discusses possible numerical methods to achieve the optimal recovery rate, i.e., to solve the PDE to the highest accuracy possible given only the information $({\bf f},{\bf g})$ (see (1.5)). An important consequence of the analysis in §5 is that the widely used loss function ${\cal L}_{0}$ (or ${\cal L}$) may not be an appropriate discretization of the theoretical loss ${\cal L}_{T}$. We show that the correct loss function ${\cal L}^{*}:H^{1}(\Omega)\to\mathbb{R}$ to use in collocation methods in our setting is
@@
${\cal L}^{*}(v)\!:=\!\!\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}|f({\bf x}_{i})+\Delta v({\bf x}_{i})|^{\gamma}\right]^{1/\gamma}\!\!\!\!+\left[\frac{1}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|[g-v]({\bf z}_{i})-[g-v]({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{2}}\right]^{1/2}\!\!\!+\left[\frac{1}{\overline{m}}\sum_{j=1}^{\overline{m}}|g({\bf z}_{j})-v({\bf z}_{j})|^{2}\right]^{1/2},$ (1.9)
@@
where $\gamma$ is the smallest number (if this is possible) so that $L_{\gamma}(\Omega)$ embeds in $H^{-1}(\Omega)$. For example, $\gamma=\frac{2d}{d+2}$ in the case $d\geq 3$. The case $d=2$ is more complicated and we will discuss it later in the paper.

### v2

@@ Abstract
We provide an a priori analysis of a certain class of numerical methods, commonly referred to as collocation methods, for solving elliptic boundary value problems. They begin with information in the form of point values of the right side $f$ of such equations and point values of the boundary function $g$ and utilize only this information to numerically approximate the solution $u$ of the Partial Differential Equation. For such a method to provide an approximation to $u$ with guaranteed error bounds, additional assumptions on $f$ and $g$, called model class assumptions, are needed. We determine the best error (in the energy norm) of approximating $u$, in terms of the total number of point samples, under all Besov class model assumptions for the right hand side and boundary data.
We then turn to the study of numerical procedures and analyze whether a proposed numerical procedure (nearly) achieves the optimal recovery error. In particular, we analyze numerical methods which generate the numerical approximation to $u$ by minimizing specified data driven loss functions over a set $\Sigma$ which is either a finite dimensional linear space, or more generally, a finite dimensional manifold. We show that the success of such a procedure depends critically on choosing a data driven loss function that is consistent with the PDE and provides sharp error control. Based on this analysis a loss function ${\cal L}^{*}$ is proposed.
We also address the recent methods of Physics Informed Neural Networks. We prove that minimization of the new loss ${\cal L}^{*}$ over restricted neural network spaces $\Sigma$ provides an optimal recovery of the solution $u$, provided that the optimization problem can be numerically executed and $\Sigma$ has sufficient approximation capabilities. We also analyze variants of ${\cal L}^{*}$ which are more practical for implementation. Finally, numerical examples illustrating the benefits of the proposed loss functions are given.
@@ 1 Introduction
where $\partial\Omega$ is the boundary of $\Omega$. In order to prove the existence and uniqueness of a solution to (1), one needs to impose conditions on $f,g,a$, and $\Omega$. A standard set of assumptions that guarantee the existence and uniqueness of a solution to (1) is the following: A1: $\Omega$ is a Lipschitz (graph) domain, i.e., $\Omega$ is an open connected set in $\mathbb{R}^{d}$ with a Lipschitz boundary in the sense described in [67]; A2: $f\in H^{-1}(\Omega)$; A3: $g\in H^{1/2}(\partial\Omega)$; A4: the diffusion coefficient $a$ satisfies the coercivity condition
$0<M_{0}\leq a({\bf x})\leq M,\quad{\bf x}\in\Omega,$ (1.2)
for some constants $M_{0},M>0$. Here the spaces $H^{s}(\Omega)$, $s\in\mathbb{R}$, are the Sobolev spaces of order $s$ in $L_{2}(\Omega)$.
@@
where the constants $c,C$ depend on $\Omega$, $M_{0}$ and $M$  [^1] Throughout this paper, we use the notation $c,C,c_{1},C_{1}$, etc, to denote constants. Unless they are absolute constants, we indicate the quantities on which they depend when they occur. The value of these constants can change at each occurrence. We also use the notation $A\asymp B$ to denote that $c_{1}A\leq B\leq c_{2}A$ with $c_{1},c_{2}>0$.. Note that assumptions A2-A3 could be replaced with other smoothness assumptions on $f$ and $g$, respectively, which would result in a theory similar to the one described in this paper.
In order to make our presentation as clear as possible, we want to avoid the technicalities of handling general domains and general diffusion coefficients $a$. Accordingly, we restrict our presentation to the case $\Omega=(0,1)^{d}$ and $a\equiv 1$. This will allow us to concentrate on the novelties of PINNs and alternative collocation methods. We will comment on the extension of our results to more general domains $\Omega$ in §11.
Note that because of (1.3), we have that for every $v\in H^{1}(\Omega)$
$\|u-v\|_{H^{1}(\Omega)}\asymp\|f+\Delta v\|_{H^{-1}(\Omega)}+\|g-v\|_{H^{1/2}(\partial\Omega)},$ (1.4)
with absolute constants in the equivalence. Hence, the problem of finding the minimum of the function ${\cal L}_{T}$ (which we call theoretical loss function),
@@
over the whole of $H^{1}(\Omega)$ will have $u$ as its unique solution, that is,
$u=\mathop{\rm argmin}_{v\in H^{1}(\Omega)}{\cal L}_{T}(v).$ (1.6)
@@ 1.1 Numerical methods for solving (1)
A numerical method assumes that $f,g$, and $\Omega$ are known and suggests a procedure for solving (1) based on that information. The numerical procedure generates a function $\hat{u}$ which is an approximation to $u$. The efficiency of the numerical method is then evaluated in the following sense. One chooses a norm $\|\cdot\|_{X}$ in which to measure performance and then seeks to quantify how fast the error $\|u-\hat{u}\|_{X}$ between the true solution $u$ and the output $\hat{u}$ of the numerical procedure tends to zero in terms of the complexity of the numerical procedure. The classical norm $\|\cdot\|_{X}$ used to measure error is the energy norm which corresponds to choosing $X=H^{1}(\Omega)$. Other function norms sometimes used to measure error correspond to $X=L_{p}(\Omega)$, $1\leq p\leq\infty$. We restrict our analysis of convergence and rates of convergence to the case $X=H^{1}(\Omega)$ in going further. We remark that our approach can in principle be applied to other error norms as well. In fact, the case of $X=H^{1/2}(\Omega)$ has been considered in a similar way in [78].
An important question is how one can compare the performance of different numerical methods for solving (1) in order to give a fair competition between all possible numerical methods. This is typically done by assigning some form of complexity to the numerical method. In the case of (A)FEMs, this complexity can be measured in terms of the number $n$ of machine operations used to compute $\hat{u}_{n}=\hat{u}$, and there are theorems that give a priori bounds for the error $\|u-\hat{u}_{n}\|_{H^{1}(\Omega)}$ under additional assumptions on the smoothness of $f$ and $g$ (see e.g. [6]). Some alternatives to measuring the complexity $n$ in terms of machine operations are the following. If the numerical approximation $\hat{u}$ to $u$ is an element from a linear $n$-dimensional space $V_{n}$, then the associated complexity is typically assigned to be the dimension $n$. This may not be directly converted to an appropriate number of machine operations because numerical stability becomes an issue. If the approximation $\hat{u}$ comes from a nonlinear manifold ${\cal M}_{n}$, then one can use the number of parameters $n$ determining the manifold as a complexity measure. However, in the latter case, one has to impose extra conditions on the manifold (or the numerical procedure) to prevent the use of space filling manifolds and thus avoid unstable numerical methods, see [10]. In the absence of such restrictions, the numerical method may require an inordinate amount of computational resources to achieve a desired accuracy even in the case when the dimension $n$ is small.
@@
A second item in collocation methods is how one numerically uses the $m$ pieces of information provided to create a numerical approximation to $u$. If $\Sigma_{n}$ is the underlying space used to create $\hat{u}$ then the accuracy of the numerical recovery will depend on $n$ and the choice of $\Sigma_{n}$. In other words, $\hat{u}$ can be viewed as $\hat{u}=\hat{u}_{n,m}$. If one fixes $m$, then one can study how fast the accuracy of approximation $\|u-\hat{u}_{n,m}\|$ tends to the optimal recovery rate as $n\to\infty$. This issue is discussed in §9. We remark that when neural network based methods such as PINNs are used to perform this numerical approximation, a critical component of the numerical algorithm becomes the optimization method used to train the neural network. This issue, which is specific to neural networks, is addressed in detail in §10.
In going further, we consider problem (1) under the following assumptions: B1: $\Omega=(0,1)^{d}$, $d\geq 2$; B2: $f\in C(\overline{\Omega})\subset H^{-1}(\Omega)$; B3: $g\in H^{1/2}(\partial\Omega)\cap C(\partial\Omega)$; B4: $a\equiv 1$.
@@ 1.2 Collocation methods and PINNs
$\hat{u}\,\in\,\mathop{\rm argmin}_{S\in\Sigma_{n}}{\cal L}(S),$ (1.8)
@@
${\cal L}(S):=\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}[\Delta S({\bf x}_{i})+f({\bf x}_{i})]^{2}\right]^{1/2}+\left[\frac{1}{\overline{m}}\sum_{i=1}^{\overline{m}}[S({\bf z}_{i})-g({\bf z}_{i})]^{2}\right]^{1/2}.$ (1.9)
Note that this is equivalent to the square root of the original PINNs loss [59]. Later in the paper we will consider weighted versions of ${\cal L}$. The appropriateness of the loss ${\cal L}$ and how $\widetilde{m},\overline{m}$, and the data sites $({\cal X},{\cal Z})$ should be chosen are important issues and are one of the focal points of this paper. We remark that an analysis of other methodologies for solving (1) using neural networks, such as the deep Ritz method [73] and finite neuron method [74], can be found in [51, 49, 66, 56, 28, 26].
The PINNs numerical procedures fall into the broad class of collocation methods for solving PDEs, i.e. numerical methods that use only values of $f$ and $g$ at specified data sites $({\cal X},{\cal Z})$. Such methods were studied in the last century (see e.g. [2, 41, 38]) but became less popular with the advent of FEMs. The novelty of PINNs is to use neural networks instead of polynomials or splines to build the approximation $\hat{u}$ from the given data.
@@ 1.3 Overview of this paper
The idea of using neural networks to solve PDEs goes back to the last century [25, 47], and has recently become extremely popular with the introduction of PINNs [59]. Despite the increasing usage of PINNs for numerically solving a wide range of PDEs (see for example, [43, 11, 7, 52] and the references therein), a satisfactory analysis of the convergence and performance of these methods has not yet been given. We refer to [19] and the references therein for an overview of the existing approaches. Some partial progress towards this goal has recently been made, see for instance [31, 54, 63, 64, 78].
@@
For example, in [63], convergence is proved in the $C(\overline{\Omega})$-norm for a modified physics-informed neural network for elliptic PDEs under the assumption that the right hand side $f$ and boundary data $g$ are Hölder continuous. Under the additional assumption that the outputs of the neural network satisfy the boundary conditions exactly, convergence is obtained in $H^{1}(\Omega)$. In this situation, the boundary values are not enforced via the loss function and must instead be implemented through the neural network architecture. We remark that although methods for exactly enforcing boundary conditions with neural networks have been proposed in [47, 48, 4], this approach is unable to rigorously handle arbitrary boundary conditions on general domains.
@@
Further, in [64] and [78], an abstract framework for analyzing methods based upon residual minimization, such as PINNs, is presented. This framework has been used in a variety of follow up works to analyze PINNs (see [3, 68, 27] for instance).
An investigation of the PINNs method for solving a variety of equations, obtaining upper bounds on the rate of convergence with respect to the number of collocation points, has been obtained in [54]. In [50], a mini-max analysis for various neural network PDE formulations (with built in zero boundary conditions) is given for the Poisson equation with zero boundary values under the assumption that noisy values of the right hand side of the PDE are measured at random points.
Our work differs from the previous approaches in the following important ways:
  • we consider general Sobolev and Besov smoothness assumptions on the RHS and boundary data (instead of just the Hilbert space setting as is typical in prior work) and consider error in the energy norm $H^{1}(\Omega)$ (in constrast to the weaker norms often considered in prior work); Besov and Sobolev regularity is the standard staple of formulating regularity theorems for PDEs;
  • we obtain matching upper and lower bounds on the theoretically optimal convergence rate in terms of the number of collocations points;
  • we show that a properly designed loss function can be used to obtain an a posteriori error estimate on the solution error for the elliptic PDE;
  • based upon our analysis, we propose new loss functions, including a loss function which correctly balances the boundary and domain error in the original PINNs formulation, and numerically demonstrate their improved performance on problems with (relatively) low regularity.
In particular, our analysis concentrates on the following questions:
@@
Q1: Given the model class assumptions on $f$ and $g$, which query sites should be used and how large would we have to take $\widetilde{m}$ and $\overline{m}$ to be able to reach a target accuracy $\varepsilon$ for the error $\|u-\hat{u}\|_{H^{1}(\Omega)}$? Q2: Given a budget of $m$ queries and given model class assumption on $f$ and $g$, what is the smallest error that can be achieved in the recovery of $u$? This is called the optimal recovery accuracy. Q3: If we use a NN space $\Sigma_{n}$ with $n$ parameters to build the approximation $\hat{u}$ to $u$, then how large do we need to choose $n$ to achieve near optimal accuracy? Q4: If we use minimization of a data driven loss function to generate the numerical approximation $\hat{u}$, then what form should the loss function take so that small values of this loss function guarantee small values of the solution approximation error, up to the optimal recovery rate ?
@@
We answer the above questions for all Besov space model classes and thereby establish such an a priori analysis of collocation methods and in particular PINNs. In contrast to previous works, we consider the error in the $H^{1}(\Omega)$-norm, and treat the situation where the boundary values are enforced through the loss function. We assume that $f\in{\cal F}$, $g\in{\cal G}$, where ${\cal F}$ and ${\cal G}$ are compact subsets of the space of continuous functions on the domains $\overline{\Omega}$ and $\partial\Omega$, respectively. Such compact sets are typically described by smoothness conditions. In this paper, we measure smoothness by membership in Besov classes. For this reason, we begin in §2 with the definition of Besov spaces and the properties we will need in the sequel. We also discuss there how well functions in Besov classes can be approximated (in various norms) by piecewise polynomials. All of the results in that section are known and are therefore presented in a summary form. However, for completeness of the paper, and for the specific presentation of these results, we give proofs of the results we need in the appendix.
@@
We use membership in Besov spaces $B_{q}^{s}(L_{p}(\Omega))$ as the smoothness we impose on $f$ and $g$. Once such assumptions are placed on $f$ and $g$, we answer questions Q1-Q2 in §3. Questions of this type are commonly referred to as optimal recovery (OR) of a function $u$ from data. Since our data is given by point samples, this is also known as optimal sampling. Our general setting is not usually addressed in the OR literature because our data is not point evaluations of the target function $u$, but rather of $f$ and $g$. Nevertheless, we consider the content of §3 to be for the most part known in the sense that we are simply piecing together various known results and techniques such as those in [13, 14, 45, 58]. The paper [69], which studies optimal sampling with respect to the negative Sobolev norm $H^{-1}(\Omega)$, comes closest to presenting the optimal recovery results in the form we need. However, note that the cases $10<p,q<$ 1 for all $d$, or $p=1$, $0<q\leq\infty$, $d=2$ are not covered in [69]. Notably, we observe that the optimal sampling rates for Besov spaces in $H^{-1}(\Omega)$ coincide with the optimal sampling rates in a certain space $L_{\gamma}(\Omega)$, determined by the Sobolev embedding (see Theorem 3.1 (iv)).
@@
The results of §3 do not give a numerical method for finding $\hat{u}$ from the given data that provides the error bound $\|u-\hat{u}\|_{X}\leq\varepsilon$. They simply establish the theoretical benchmark for the optimal accuracy of any numerical procedure for the recovery of $u$ based on the point samples of $f$ and $g$, see the discussion at the beginning of §4. Our analysis there does not involve NNs or any other proposed collocation method. NNs and PINNs only enter the picture when one wants to transform the theoretical analysis into a numerical procedure that utilizes NNs to find an approximant to $u$. We then further discuss in §4 the use of classical numerical methods such as Finite Element Methods (FEMs) for optimal recovery.
@@
In §5, we turn to the question of using optimization to solve the PDE. We observe that minimizing the theoretical loss ${\cal L}_{T}$ over a space $\Sigma_{n}$ will approximate well $u$, provided $\Sigma_{n}$ has sufficient approximation power. However, this is not a numerical algorithm per se because it does not provide a numerical recipe for evaluating ${\cal L}_{T}(S)$ when $S\in\Sigma_{n}$.
@@
The remainder of this paper discusses possible numerical methods to achieve the optimal recovery rate, i.e., to solve the PDE to the highest accuracy possible given only the information $({\bf f},{\bf g})$ (see (1.7)). An important consequence of the analysis in §5 is that the widely used loss function ${\cal L}$ may not be an appropriate discretization of the theoretical loss ${\cal L}_{T}$. We introduce the loss function ${\cal L}^{*}:H^{1}(\Omega)\to\mathbb{R}$ given by
@@
${\cal L}^{*}(v)\!:=\!\!\left[\frac{1}{\widetilde{m}}\sum_{i=1}^{\widetilde{m}}|f({\bf x}_{i})+\Delta v({\bf x}_{i})|^{\gamma}\right]^{1/\gamma}\!\!\!\!+\left[\frac{1}{\overline{m}^{2}}\sum_{\begin{subarray}{c}i,j=1\\ i\neq j\end{subarray}}^{\overline{m}}\frac{|[g-v]({\bf z}_{i})-[g-v]({\bf z}_{j})|^{2}}{|{\bf z}_{i}-{\bf z}_{j}|^{d}}\right]^{1/2}\!\!\!+\left[\frac{1}{\overline{m}}\sum_{j=1}^{\overline{m}}|g({\bf z}_{j})-v({\bf z}_{j})|^{2}\right]^{1/2},$ (1.10)
@@
and show that it correctly discretizes an upper bound on the theoretical loss ${\cal L}_{T}$. Here $\gamma$ is the smallest number (if this is possible) so that $L_{\gamma}(\Omega)$ embeds in $H^{-1}(\Omega)$. For example, $\gamma=\frac{2d}{d+2}$ in the case $d\geq 3$. The case $d=2$ is more complicated and we will discuss it later in the paper.